\begin{thebibliography}{99}
\bibitem{Albin} P. Albin, \emph{A renormalized index theorem for some complete asymptotically regular metrics: the Gauss-Bonnet theorem}, Adv. in Maths,, 213(1):1--52, 2007.
\bibitem{BLS} K. Bohlen, J.-M. Lescure, E. Schrohe, \emph{The Atiyah-Patodi-Singer index formula for pseudodifferential operators on Lie structures}, in preparation. 
\bibitem{BohlenSchrohe} K. Bohlen, E. Schrohe, \emph{Getzler rescaling via adiabatic deformation and a renormalized index formula}, J. Math. Pures Appl. 120 (2018): 220-252.
\bibitem{DH} S. Dave, S. Haller, \emph{Graded hypoellipticity of BGG sequences}, Annals of Global Analysis and Geometry 62.4 (2022): 721-789.
\bibitem{DS} C. Debord, G. Skandalis, \emph{Adiabatic groupoid, crossed product by $\mathbb{R}_{+}^{\ast}$ and pseudodifferential calculus}, Adv. Math. 257 (2014), pp. 66--91.
\bibitem{DebordSkandalis} C. Debord, G. Skandalis, \emph{Blow-up constructions for Lie groupoids and a Boutet de Monvel type calculus}, M\"unster Journal of Mathematics 14.1 (2021).
\bibitem{vEY} E. van Erp, R. Yuncken, \emph{A groupoid approach to pseudodifferential calculi}, Journal f\"ur die reine und angewandte Mathematik (Crelles Journal) 2019.756 (2019): 151-182.
\bibitem{Grad} \v{Z}. Grad, \emph{Fundamentals of Lie categories and Yang-Mills theory for multiplicative Ehresmann connections} arXiv preprint arXiv:2507.08220 (2025).
\bibitem{Greiner} P. Greiner, \emph{An asymptotic expansion for the heat equation}, Arch. Rational Mech. Anal. 41, 163--218 (1971).
\bibitem{Heitsch} J. L. Heitsch, \emph{Bismut super-connections and the Chern characters for Dirac operators on foliated manifolds}, K-Theory, 9:507--528, 1995.
\bibitem{SLLS} A. R. H. S. Sadegh, S. Liu, Y. Loizides, J. Sanchez, \emph{A fixed-point formula for Dirac operators on Lie groupoids}, Journal of Functional Analysis, 287(11), 2024: 110624.
\bibitem{LR} N. P. Landsman, B. Ramazan, \emph{Quantization of Poisson algebras associated to Lie algebroids}, Contemporary Mathematics 282 (2001).
\bibitem{LV} J.-M. Lescure, S. Vassout, \emph{On evolution equations for Lie groupoids}, arXiv preprint, arXiv:2010.00227 (2020).
%\bibitem{N} V. Nistor, \emph{Pseudodifferential operators on non-compact manifolds and analysis on polyhedral domains}, Workshop on Spectral Geometry of Manifolds with Boundary and Decomposition of Manifolds, Roskilde University, 2005, S. 307-328.
\bibitem{McDonald} E. McDonald, \emph{$L_p$ estimates in the Androulidakis-Mohsen-Yuncken calculus}, arXiv preprint arXiv:2410.13701 (2024).
\bibitem{Melrose} R. Melrose, \emph{The Atiyah-Patodi-singer index theorem}, AK Peters/CRC Press, 1993.
\bibitem{Monthubert} B. Monthubert, \emph{Pseudodifferential calculus on manifolds with corners and groupoids}, Proc. Amer. Math. Soc. 127.10 (1999), pp. 2871-2881.
\bibitem{NWX} \emph{Pseudodifferential operators on differential groupoids}, Pacific journal of mathematics 189.1 (1999): 117-152.
\bibitem{Piriou} A. Piriou, \emph{Une classe d'op\'{e}rateurs pseudo-diff\'{e}rentiels du type de Volterra}, Ann. Inst. Fourier 20, 77--94 (1970).
\bibitem{PongeMika} R. Ponge, H. Mikayelyan, \emph{On the asymptotic completeness of the Volterra calculus}, Journal d'Analyse Mathematique, 94(1), 249-263, 2004.
\bibitem{Skandalis_course_2015} G. Skandalis, \emph{{C}${}^*$-alg\`ebres, {A}lg\`ebres de {V}on {N}eumann, {E}xemples}, Cours de M2, 2015, \href{https://webusers.imj-prg.fr/~georges.skandalis/poly2015.pdf}{https://webusers.imj-prg.fr/$\sim$georges.skandalis/poly2015.pdf}.
\bibitem{So} B. K. So, \emph{Exponential coordinates and regularity of groupoid heat kernels}, Central European Journal of Mathematics, Volume 12, pages 284--297, (2014).
%\bibitem{SoThesis} B. K. So, \emph{Pseudo-differential operators, heat calculus and index theory of groupoids satisfying the Lauter-Nistor condition}, arXiv preprint, arXiv:1006.5623 (2010).
\bibitem{Trotter} H. F. Trotter, \emph{Approximation and perturbation of semigroups}, Linear Operators and Approximation II: Proceedings of the Conference held at the Oberwolfach Mathematical Research Institute, Black Forest, March 30--April 6, 1974 (pp. 3-21), Basel: Birkh\"auser Basel.
\bibitem{Vassout} S.~Vassout, \emph{Unbounded pseudodifferential calculus on {L}ie groupoids},
{\em J. Funct. Anal.}, 236(1):161--200, 2006.
%\bibitem{vEY} E. Van Erp, R. Yuncken, \emph{A groupoid approach to pseudodifferential calculi}, Journal für die reine und angewandte Mathematik (Crelles Journal) 2019.756 (2019): 151-182.
\end{thebibliography}

\documentclass[11pt, reqno]{amsart}
\usepackage{amsmath, amsthm, amscd, amsfonts, amssymb, graphicx, color, stmaryrd, mathrsfs} 
\usepackage[all,cmtip]{xy}	
\usepackage[bookmarksnumbered, colorlinks, plainpages]{hyperref}
\usepackage{float}	
\usepackage{tikz-cd}

\tikzset{%
    symbol/.style={%
        ,draw=none
        ,every to/.append style={%
            edge node={node [sloped, allow upside down, auto=false]{$#1$}}}
    }
}

\def\presuper#1#2{\mathop{}\mathopen{\vphantom{#2}}^{#1}\kern-\scriptspace#2}

\newcommand{\Nn}{\mathbb{N}}
\newcommand{\Rr}{\mathbb{R}}
\newcommand{\Cc}{\mathbb{C}}
\newcommand{\Zz}{\mathbb{Z}}

\newcommand{\Dom}{\mathrm{dom}}	
\newcommand{\F}{\mathcal{F}}
\newcommand{\K}{\mathcal{K}}	
\renewcommand{\L}{\mathcal{L}}	
\newcommand{\E}{\mathcal{E}}	
\newcommand{\G}{\mathcal{G}}	
\newcommand{\A}{\mathcal{A}} 	
\newcommand{\U}{\mathfrak{A}}  	
	
\newcommand{\End}{\mathrm{End}}	

\newcommand{\Hom}{\mathrm{Hom}}

\newcommand{\Op}{\mathrm{Op}}

\newcommand{\Ff}{\F_{\mathrm{fib}}}	

\renewcommand{\L}{\mathcal{L}} 

\newcommand{\Diff}{\mathrm{Diff}}	
\newcommand{\supp}{\mathrm{supp}}	

\newcommand{\Id}{\operatorname{Id}}	

\newcommand{\Aut}{\mathrm{Aut}}

\newcommand{\scal}[2]{\langle #1, #2 \rangle}	

\newcommand{\dtimes}{\bar{\times}}
\newcommand{\pa}{\mathrm{pa}}
\newcommand{\Exp}{\mathrm{Exp}}

\newcommand{\iso}{\xrightarrow{\sim}}	

\newtheorem{Thm}{Theorem}[section]
\newtheorem{Lem}[Thm]{Lemma}
\newtheorem{Prop}[Thm]{Proposition}

\theoremstyle{definition}
\newtheorem{Def}[Thm]{Definition}

\newtheorem{Rem}[Thm]{Remark}

\begin{document}
\setcounter{page}{1}


\title[Volterra Calculus]{A Volterra Calculus for Lie Groupoids}


\author[Karsten Bohlen]{Karsten Bohlen}



\subjclass[2000]{Primary 47J35; Secondary 58H05, 35K67.}

\keywords{groupoid, Volterra calculus, heat kernel}



\begin{abstract}
A pseudodifferential Volterra calculus for inverting parabolic differential equations on Lie groupoids is introduced. This enables the study of fundamental solutions of various cases of heat flows on singular manifolds with corners with non-resonant boundary indicial symbols, such as the $b$-manifolds, as well as other geometric bisection covariant heat flows. We also establish the short time asymptotic expansion for the heat kernel of a positive, elliptic differential operator on a Lie groupoid that acts on suitable Sobolev Hilbert modules and is positive definite with respect to the appropriate $L^2$ inner product.
\end{abstract} \maketitle 


\section{Introduction}


The Volterra calculus was introduced by P. Greiner \cite{Greiner} and A. Piriou \cite{Piriou} in order to study the fundamental solutions of parabolic equations, such as the heat equation. This construction has been extended to various situations, cf. \cite{PongeMika}. For instance, R. Melrose extended the calculus to investigate the $b$-differential operators on compact manifolds with boundary and used it to give a heat kernel proof of the Atiyah-Patodi-Singer index theorem \cite{Melrose}. The (small) $b$-calculus of pseudodifferential operators of R. Melrose was extended to $b$-groupoids by B. Monthubert \cite{Monthubert} and there followed many similar groupoid constructions for other types of singular manifolds. We refer to \cite{DebordSkandalis} for a unified approach to blowup constructions on Lie groupoids. A pseudodifferential calculus for Lie groupoids has been described by V. Nistor, A. Weinstein and P. Xu \cite{NWX}. Our aim here is to introduce a Volterra calculus on a given Lie groupoid that enhances the existing pseudodifferential calculus to deal with parabolic equations, stated with respect to elliptic right invariant differential operators. In particular this enables a study of the heat semigroup defined on a Lie groupoid and to establish its short time asymptotic expansion. It should be noted that there are possible logarithmic terms that may arise in the asymptotic expansion, which can be local in origin or arise from the boundary resonances of the indicial symbols on a manifold with corners, especially in more singular cases. Secondly, there is a constraint involving possible limited transverse regularity. The issue of the regularity of the heat semigroup on a Lie groupoid has been studied e.g. by B. K. So \cite{So} and in the case of regular folations by J. L. Heitsch \cite{Heitsch}. With the exception of a few cases such as that of $b$-type \cite{Melrose} and edge type singularities \cite{Albin} on manifolds with boundary, the holonomy groupoid of regular foliations \cite{Heitsch}, as well as asymptotically Euclidean boundary groupoids \cite{So}, it is generally not known if the heat kernel can always be smoothed out in the transverse direction. The third issue that one encounters in the construction of pseudodifferential calculi on Lie groupoids is that the inverse operator of invertible, elliptic elements of the calculus may not be contained in the calculus due to the required proper support conditions. In the case of the Volterra calculus a suitable choice of sufficiently well-behaved smoothing ideal ensures the correct smoothing behavior outside of the diagonal. Recently, in work due to E. van Erp and R. Yuncken \cite{vEY} a groupoid approach to pseudodifferential operators is described; compare also the work of Debord-Skandalis, cf. \cite{DS}. The idea is to access an element in the pseudodifferential calculus on a manifold by lifting it to a family of operators on the tangent groupoid that glues, in a precise sense, the operator to its principal symbol. This vantage point allows also an elegant treatment of extensions to the standard pseudodifferential calculus that involve a filtration of the tangent space or Lie algebroid, such as the anisotropic operators that constitute the Volterra heat calculus. Such a calculus may have applications to the study of certain bisection covariant longitudinal heat flows, cf. \cite{Grad}. We should remark however that we can not simply recover the Volterra calculus described herein as a particular instance of E. van Erp and R. Yuncken, due in large part to the aforementioned difficulties. However, one is able to obtain a heat calculus this way in the case of Lie groupoids over a compact base that arise from Lie group actions, cf. \cite{SLLS}.

\subsection*{Overview}

We fix a compact manifold with corners $M$ of dimension $d \in \Nn$ and a Lie groupoid $\G \rightrightarrows \G_0 = M$ with a vector bundle $E \to M$. 
On the compact manifold with corners $M$, we consider an induced filtration of the Lie algebra of vector fields on $M$ that is of a parabolic signature. 

We denote by $\overline{r}, \overline{s} \colon \G \dtimes \Rr \rightrightarrows M \times \{0\} = M$ the groupoid formed as product with the additive group $(\Rr, +)$. The calculus is defined as
\begin{align}
\widehat{\Psi}_{ai}^{m}(\G \dtimes \Rr, \overline{r}^{\ast} E) &:= \Psi_{ai}^{m}(\G \dtimes \Rr, \overline{r}^{\ast} E) + \Psi_{\G \dtimes \Rr, ai}^{-\infty}(\overline{r}^{\ast} E), \label{LargeVolterraCalculus}
\end{align} 
i.e. the linear span formed out of the quantization of anisotropic symbols of order $m$, that yield compactly supported Schwartz kernels, with the addition of residual operators that belong to the class of the Vassout ideal $\Psi_{\G \dtimes \Rr}^{-\infty}$ (cf. \cite{Vassout}) on the product groupoid, with the additional Volterra or causality property. Elements of the Vassout ideal are constructed out of Hilbert $C^{\ast}$-modules and by definition reside inside the reduced $C^{\ast}$-algebra, $C_r^{\ast}(\G \dtimes \Rr)$. It can be shown however that elements of this ideal are included in the class of kernels which are transversely continuous and smooth along every orbit of the groupoid, by an application of the right-regular representation, i.e.
\[
\Psi_{\G \dtimes \Rr}^{-\infty} \subset C_{\mathrm{orb}}^{\infty, 0}(\G \dtimes \Rr) \cap C_r^{\ast}(\G \dtimes \Rr), 
\] 
cf. \cite[Thm. 10]{LV}. The first component in the definition of the large calculus is formed out of symbols which are smooth along the $s$-fibers, of type $C^{\infty, 0}$. Altogether, therefore we have the inclusion
\begin{align}
\widehat{\Psi}_{ai}^m(\G \dtimes \Rr, \overline{r}^{\ast} E) &\subset C^{\infty, 0}(\G \dtimes \Rr, \overline{r}^{\ast} E) \cap C_r^{\ast}(\G \dtimes \Rr, \overline{r}^{\ast} E). \label{LongContInclusion} 
\end{align}

A further refinement of this inclusion is possible when we utilize the Volterra property and the Paley-Wiener theorem. Then we get an inclusion into the so-called Paley-Wiener algebra intersected with the class of $C^{\infty, 0}$ kernels, by virtue of an application of the fiberwise Fourier transform. We return to this point at a later stage. Moreover, we will exhibit for each pseudodifferential operator $T \in \widehat{\Psi}_{ai}^{m}(\G \dtimes \Rr, \overline{r}^{\ast} E)$ a $1$-parameter family $\{T_{\hbar}\}_{\hbar > 0}$ with $T_{1} = T$. To that end we construct a suitable deformation groupoid for our filtration of parabolic signature which constitutes the proper weighted generalization of the tangent groupoid, cf. \cite{vEY}. This groupoid with respect to the filtration $H$ is given as a set as 
\begin{align}
(\G \dtimes \Rr)^{\pa} &= [(\G \dtimes \Rr) \times \Rr_{\hbar}^{\ast}] \cup [(\G \dtimes \Rr)_{H} \times \{0\}_{\hbar}] \rightrightarrows M \times \Rr_{\hbar}. \label{DeformationGroupoid}
\end{align}

We utilize throughout this work the generalized exponential map to exhibit a localizer and a so-called standard system of coordinates wherein which we can express elements of the small compactly supported calculus, while elements of the Vassout ideal are understood via the right-regular representation. The important properties of the calculus, such as the closedness with respect to compositions and the parametrix construction, can then be verified in the standard coordinates. 




\section{The heat semigroup}


We recall the construction of the heat semigroup on a Lie groupoid, cf. \cite{BLS}. Let $\E := C_r^{\ast}(\G, r^{\ast} E)$ be the $C_r^{\ast}(\G)$ module. We also recall the definition of the Sobolev Hilbert modules, cf. \cite{Skandalis_course_2015}. 
Denote the calculus of pseudodifferential operators with compactly supported Schwartz kernels by $\Psi_{\G, c}^{\ast}$. Fix the \emph{generator} $P \in \Psi_{c, \G}^{1}$ and define the scale of Sobolev modules as the $C_r^{\ast}(\G)$-modules $H^t := \overline{C_c^{\infty}(\G)}^{\scal{\cdot}{\cdot}_t}, \ t \in \Rr$, where 
\[
\scal{f}{g}_t = \scal{(\Id + P^2)^t f}{g} \in C_r^{\ast}(\G), 
\]

and $\scal{a}{b} = a^{\ast} b$ is the inner product on the right. Given $Q \in \Psi_{\G, c}^m$ we have that $Q \in \L(H^t, H^{t-m})$, where the latter class of so-called \emph{morphisms} are the $C_r^{\ast}(\G)$-linear maps that admit an adjoint, or equivalently, the $C_r^{\ast}(\G)$-linear maps that have orthocomplemented graphs. Set:
\begin{align}
\|Q\|_{t\to t-m} := \|(1+P^2)^{(t-m)/2}Q(1+P^2)^{-t/2}\|_{\L(\E)}. \label{SovolevNorm}
\end{align}

Define the Vassout smoothing operators $\Psi_{\G}^{-\infty}$ (cf. \cite{Vassout}) via
\begin{align}
\E^{\infty} &:= \bigcap_k \Dom(P^k), \ \Psi_{\G}^{-\infty} := \E^{\infty} \cap (\E^{\infty})^{\ast}. \label{VassoutIdeal} 
\end{align}
Note that $\Psi_{\G, c}^{0}$ constitutes a subalgebra of the multiplier algebra $M(C_r^{\ast}(\G))$ and $\Psi_{\G, c}^{<0}$ is contained in $C_r^{\ast}(\G)$. In addition,
\[
H^t = \overline{\Psi_{\G, c}^{<-t}}^{\|\cdot\|_t}, \ \E^{\infty} = \bigcap_{t} H^t \subset C_r^{\ast}(\G). 
\]

The compact morphisms $\E \to \F$ between Hilbert $A$-modules are defined as the linear subspace spanned by the usual rank one projections; the embedding $H^t \hookrightarrow H^{t+m}$ is always compact in that sense, for each $m > 0$. A densely defined operator $T \colon \E \to \E$ for a Hilbert $A$-module $\E$, with densely defined adjoint is regular if its graph $G(T)$ is orthocomplemented as an $A$-submodule of $\E \oplus \E$. The topology of strict continuity on $\L(\E)$ is the weakest topology generated by the seminorms $T \mapsto \|T x\|, x \in \E$ and $T \mapsto \|T^{\ast} y\|, y \in \E$. It can be checked that $\L(\E)$ specified in this way is a $C^{\ast}$-algebra. If we view a $C^{\ast}$-algebra $A$ as a right Hilbert $A$-module with $\scal{a}{b} = a^{\ast} b$, then the multiplier algebra $M(A)$ identifies with the morphisms $\L(A)$. 

Consider an element $Q \in \Psi_{\G, c}^m(r^{\ast} E)$ for $m > 0$ that is self-adjoint, elliptic as an unbounded operator $\E \to \E$ and non-negative with respect to the canonical $L^2 = H^0$ inner product. Let $\pi \colon C_0(\Rr) \to \L(\E)$ be a nondegenerate representation such that we can find an extension $\widetilde{\pi}$ from $C(\Rr)$ (i.e. the regular operators on $C_0(\Rr)$) defined via $C_0(\Rr) \otimes_{\pi} \E \cong \E$ through $\widetilde{\pi}(f) = f \otimes_{\pi} \Id$ and with the property $\widetilde{\pi}(\Id_{\Rr}) = Q$. With respect to the identifications $C_b(\Rr) \cong M(C_0(\Rr))$ and $\L(\E) \cong M(\K(\E))$, we obtain a strictly continuous homomorphism by restriction
\[
\overline{\pi} = \widetilde{\pi}_{|C_b(\Rr)} \colon C_b(\Rr) \to \L(\E). 
\]

Then $\pi$ is uniquely determined by this condition, cf. \cite{LV}. 

\begin{Def}
Denote by $\Gamma$ the curve parametrized by $\Rr_{+} \ni t \mapsto -1 + t(1 \pm i)$. Define $f_{\lambda} \in C_0(\Rr)$ via $f_{\lambda}(x) = \alpha(x) (x - \lambda)^{-1}, \ \lambda \in \Gamma$. Here $\alpha$ is the continuous non-decreasing function defined by
\[
\alpha(x) = \begin{cases} 0, \ x < -\frac{1}{2}, \\ 1, \ x \geq -\frac{1}{4} \end{cases}.
\]

Then $\pi(f_{\lambda}) = (Q - \lambda)^{-1}$ and the \emph{heat kernel} is defined by the Dunford integral
\[
\exp(-tQ) := \frac{1}{2\pi i} \int_{\Gamma} e^{-t \lambda} (Q - \lambda)^{-1}\,d\lambda. 
\]

\label{Def:HeatSemigroup}
\end{Def}

\begin{Prop}
Let $Q \in \Psi_{\G,c}^m(r^*E)$ with $m > 0$ be symmetric on $C_c^\infty(\G ,r^*E)$, $\G$-elliptic, and nonnegative. Then its closure on $H^0$ is regular selfadjoint with $\mathrm{spec}(Q) \subset [0,\infty)$. For a given $k \in \Nn$ and any $u_0 \in H^k$, the initial value problem 
\[
(\partial_t + Q) u = 0, \ u(0) = u_0
\]

has a unique solution in $\bigcap_{0 \leq j \leq k} C^j([0, \infty), H^{k-j})$, given by $u(t) = \exp(-t Q)$. In addition, the heat kernel $\exp(-t Q)$ furnishes an $H^0 = L^2$ contractive semigroup that is contained in $\Psi_{\G}^{-\infty}(r^{\ast} E)$.
\label{Prop:HeatSemigroup}
\end{Prop}

\begin{proof}
We observe that $Q$ is closable and its closure is a regular selfadjoint unbounded morphism on $\E$ to which we apply the continuous functional calculus. We have $Q \in \L(H^{s + m}, H^{s})$ for the scale of Sobolev modules $\{H^s\}$ where $H^s$ is the domain of $(\Id + Q)^{\frac{s}{m}}$ with equivalent norm, by ellipticity. Denoting by $\pi : C_0(\Rr) \to \L(\E)$ the functional calculus associated with $Q$, we get $\pi(f_\lambda) = (Q-\lambda)^{-1}$ and $\|(Q-\lambda)^{-1}\|_{\E} \le \|f_\lambda\|_{\infty}\le c(1+|\Im \lambda|)^{-1}$ for some constant $c$. It follows that the integral above is absolutely convergent in $\L(\E)$, and standard calculations on this integral prove that $P^ke^{-tQ}$ and $e^{-tQ}P^k$ belong to $\L(\E)$ as well, for any $k\in\Nn$. By construction of the functional calculus, the map $\lambda \mapsto \pi(f_{\lambda})$ is strictly continuous, cf. \cite[Proposition 5.19]{Skandalis_course_2015}. By virtue of the identity
\[
\frac{(Q - \lambda)^{-1} - (Q - \mu)^{-1}}{\lambda - \mu} = \frac{(Q - \lambda)^{-1} (\mu - \lambda) (Q - \mu)^{-1}}{\lambda - \mu} = -(Q - \lambda)^{-1} (Q - \mu)^{-1} 
\]

we obtain as $\lambda \to \mu$ that the left hand side converges to $-(Q - \lambda)^{-2}$. Hence $\Cc \setminus \mathrm{spec}(Q) \ni \lambda \mapsto (Q - \lambda)^{-1}$ is holomorphic as a map $\Cc \setminus \mathrm{spec}(Q) \to \L(H^s)$. By the previous arguments we have $\exp(-t Q) \in \L(H^k)$ and $t \mapsto \exp(-t Q) f \in C(\Rr, H^{k})$ for any $f \in H^k$. Since $\frac{1}{t} (\exp(-t \lambda) - 1) \to \lambda$ as $t \to 0$ uniformly on compact subsets of $\Rr$, we get by \cite[Appendix]{Skandalis_course_2015} that $\frac{1}{t} (\exp(-t Q) - 1)$ converges to $Q$ strongly meaning
\[
\left\|\frac{1}{t} (\exp(-t Q) f - f) - Q f\right\|_{\E} \to 0, \ \text{as} \ t \to 0, \ \text{for all} \ f \in H^1.
\]

Hence $t \mapsto \exp(-t Q) f \in C^1(\Rr, \E) \cap C^0(\Rr, H^1)$ for all $f \in H^1$ and $\partial_t \exp(-t Q) f = -Q \exp(-t Q) f$ for all $f \in H^1$ and all $t \in \Rr$. Finally, we need to show that the heat kernel furnishes a semigroup and the $H^0 = L^2$ contractivity. Let $t > 0, s > 0$ and denote by $\Gamma'$ a curve that is shifted to the right slightly compared to $\Gamma$, then by Cauchy's theorem
\begin{align*}
e^{-Q t} e^{-Q s} &= (2 \pi i)^{-2} \int_{\Gamma} \int_{\Gamma'} e^{-\lambda t} (Q - \lambda)^{-1} e^{-\mu s} (Q - \mu)^{-1} \,d\mu \,d\lambda \\
&= (2\pi i)^{-2} \int_{\Gamma} \int_{\Gamma'} e^{-\lambda t - \mu s}(\mu - \lambda)^{-1} ((Q - \lambda)^{-1} - (Q - \mu)^{-1})\,d\mu \, d\lambda \\
&= (2\pi i)^{-1} \int_{\Gamma} e^{-\lambda (t + s)} (Q - \lambda)^{-1} \,d\lambda \\
&= e^{-Q (t + s)}.
\end{align*}

In the next to final line we made use of the identities
\[
\int_{\Gamma} e^{-\lambda t} (\mu - \lambda)^{-1}\,d\lambda = 0, \ \int_{\Gamma'} e^{-\mu s} (\mu - \lambda)^{-1}\,d\mu = 2\pi i e^{-\lambda s}.
\]

To show contractivity of the semigroup we can adapt a standard Hille-Yosida approximation argument to our setting, cf. e.g. \cite{Trotter}. We have 
\begin{align}
& (-\infty, 0) \subset \Cc \setminus \mathrm{spec}(Q), \label{HYRes} \\
& \|\lambda (Q + \lambda)^{-1}\| \leq 1  \label{HYBound}
\end{align}
for the regular selfadjoint operator $Q \geq 0$. Set 
\[
Q_{\lambda} := \lambda \Id - \lambda^2 (Q + \lambda)^{-1} \in \L(\E), \ \lambda > 0
\] 
and note that $Q_{\lambda} = \lambda Q (Q + \lambda)^{-1}$. Since the approximant is a bounded operator in $\L(\E)$, we obtain by virtue of the functional calculus of the $C^{\ast}$-algebra $\L(\E)$ the estimate 
\[
\|e^{-t Q_{\lambda}}\| \leq e^{-\lambda t} e^{\| \lambda^2 (Q + \lambda)^{-1}\| t} \leq e^{-\lambda t} e^{\lambda t} = 1. 
\]

For each $t \geq 0$ define the semigroup 
\[
S_t x := \lim_{\lambda \to \infty} e^{-t Q_{\lambda}} x
\] 
for $x \in \E$. One checks the uniform convergence on bounded intervals using \eqref{HYBound} which furnishes a contraction, since each $e^{-t Q_{\lambda}}$ is a contraction. One then exhibits $Q$ as a generator of $(S_t)_{t \geq 0}$ using \eqref{HYRes}, which by virtue of uniqueness of the solution to the heat equation shows that $S_t$ reproduces the heat kernel. 
\end{proof}

\section{Anisotropic Operators}




The aim is to introduce a Volterra calculus for parabolic PDE posed on a Lie groupoid. We provide a parametrix construction, so that the fundamental solutions of weakly parabolic PDE, stated with respect to the boundary tangential vector fields on the corresponding Lie algebroid, are contained within the calculus. 





The anisotropic operators arise from quantization of continuous families of Volterra symbols. These symbol classes are families of symbol spaces along the fibers of a submersion. Consider the source map $\overline{s} \colon \G \dtimes \Rr \to (\G \dtimes \Rr)_0 = M$ defined by $\overline{s}(\gamma, t) = (s(\gamma), 0)$ and typical fiber $\overline{s}^{-1}(x) = \G_x \times \Rr$. The Lie algebroid is $\A(\G \dtimes \Rr) \cong \A(\G) \oplus \Rr_M$ where $\Rr_M = M \times \Rr$ is the trivial bundle over $M$. 
A groupoid parametrization of $\overline{r}, \overline{s} \colon \G \dtimes \Rr \rightrightarrows M$, centered at $x_0 \in M$, is given by an open set $\G \dtimes \Rr \supset \Omega \cong U_{\overline{r}, \overline{s}} \times V$ for $U_{\overline{r}, \overline{s}} =: U \subset M$ open and the data $(\varphi, \psi)$ where $\varphi \colon U \to M, \ \psi \colon U \times V \to \G$ with the properties, cf. \cite{LR}
\begin{align*}
&\psi(0, 0) = x_0, \ r(\psi(u, v)) = \varphi(u) \\
&\psi(U \times \{0\}) = \psi(U \times V) \cap M,
\end{align*}

where the first two conditions imply $\varphi(u) = \psi(u, 0)$.

\begin{Def}
Let $\Cc_- := \{\tau \in \Cc : \Im \tau < 0\}$ and write the parabolic dilations as $\delta_{\lambda}(\xi, \tau) := (\lambda \xi, \lambda^2 \tau)$ for $\lambda > 0$ on the fibers of $\A^{\ast} \oplus \Rr_M$. Set $\|(\xi, \tau)\|_{ai} := (|\xi|^2 + |\tau|)^{\frac{1}{2}}$.  

Fix a groupoid parametrization $\Omega \cong U \times V$ of $\overline{r}, \overline{s} \colon \G \dtimes \Rr \rightrightarrows M$, with $U\subset M$ open and $V$ an open neighbourhood of the unit in a longitudinal chart of $\G_{x}$ (via $\psi$). For $u\in U$ write $x=\varphi(u)$ and identify the longitudinal slice
\[
V_x := \{y = \psi(u, v) \in \G_x : \ v \in V\}
\]

near the unit in $\G_x$.

\begin{enumerate}
\item For $m \in \Rr$, define $S_{ai, m}(\A^{\ast} \oplus \Rr_M)$ to be the space of families 
\[
q = (q_x)_{x \in M}, \qquad q_x \colon V_x \times (\A_x^{\ast} \times \Cc_-) \setminus \{0\} \to \Cc,
\]
such that, for each $x \in M$:
\begin{enumerate}
\item $q_x$ is $C^{\infty}$ in $\xi \in \A_x^{\ast} \setminus \{0\}$ and holomorphic in $\tau \in \Cc_-$. 

\item For all $\lambda > 0$, $q_x(y, \delta_{\lambda}(\xi, \tau)) = \lambda^m q_x(y, \xi, \tau)$. 

\item For every chart $\Omega \cong U \times V$ as above, the map
\[
u \mapsto q_{\varphi(u)}(\psi(u, \cdot), \cdot, \cdot)
\]

is continuous from $U$ into the Fr\'echet space of $C^{\infty}$ functions of $(y, \xi)$ on $V \times (\A_x^{\ast} \setminus \{0\})$ that are holomorphic in $\tau \in \Cc_-$. 
\end{enumerate}

\item A family $q = (q_x)_{x \in M}$ lies in $S_{ai, ph}^m(\A^{\ast} \oplus \Rr_M)$ if, in every chart $\Omega \cong U \times V$ and for each $x = \varphi(u) \in U$, there are fiberwise homogeneous components $q_{x, m-j} \in S_{ai, m-j}(\A^{\ast} \oplus \Rr_M)$ such that, for every $N \in \Nn$, compact $K \subset V$, multi-indices $\alpha$ in $y$, $\beta$ in $\xi$, and $k \in \Nn$,
\[
\left\|\partial_y^{\alpha} \partial_{\xi}^{\beta} \partial_{\tau}^{k} \left(q_{\varphi(u)}(\psi(u,v), \xi, \tau) - \sum_{j < N} q_{\varphi(u), m-j}(\psi(u,v), \xi, \tau) \right)\right\| \leq C \|(\xi, \tau)\|_{ai}^{m - |\beta| - 2k}.
\]

Each $q_{x,m-j}$ is $C^\infty$ in $(y,\xi)$, holomorphic in $\tau \in \Cc_-$, and parabolically homogeneous of degree $m-j$ in $(\xi,\tau)$. The constants $C$ may depend on $(\alpha,\beta,k,K)$ but are uniform over $v\in K$.
\end{enumerate}
\label{Def:anisotropicSymbol}
\end{Def}

Let $S(\A^{\ast} \oplus \Rr_M)$ (resp. $S(\A \oplus \Rr_M)$) denote fiberwise Schwartz functions on the vector bundle. We fix the convention for the fiberwise Fourier transform
\[
\Ff f(\xi, \tau) = (2 \pi)^{-(d+1)} \int_{\overline{\pi}(\xi, \tau) = \pi(\zeta, t)} e^{-i \scal{\xi}{\zeta} - i \tau t} f(\zeta, t)\,d\zeta. 
\]

\begin{Lem}[cf. \cite{Greiner}]
Let $m \in \Rr$ and $q \in S_{ai, m}(\A^{\ast} \oplus \Rr_M)$. Then there exists a unique family of tempered distributions
\begin{align*}
& g=(g_x)_{x \in M}, \ g_x(y; \xi, \tau) \in S'(\A_x^{\ast} \times \Rr),  
\end{align*}

such that 
\begin{enumerate}
\item[(i)] For each $x$, $g_x(y;\xi,\tau)=q_x(y,\xi,\tau)$ on
$(\xi,\tau)\neq (0,0)$.
\item[(ii)] For each $x$ and $\lambda>0$,
\[
g_x\left(y;\,\delta_\lambda(\xi,\tau)\right) = \lambda^{m}\,g_x(y;\,\xi,\tau)
\quad\text{in the sense of distributions in }(\xi,\tau).
\]
\item[(iii)] Writing
\[
K_x(y;\zeta,t)\;:=\;\Ff^{-1}[g_x](y;\zeta,t),
\]
we have $K_x(y;\zeta,t)=0$ for all $t<0$.
\item[(iv)] In any groupoid chart
$\Omega \cong U\times V$ and bundle trivialization of $\A^\ast$, all longitudinal seminorms in $(y,\xi)$ of $g_x$ (and of $K_x$ for $t>0$) depend continuously on $x\in U$.
\end{enumerate}
Moreover, the construction is local and chart independent, and it is functorial under
changes of longitudinal coordinates and bundle trivializations. If, in addition,
$q\in S^{m}_{ai,ph}(\A^\ast\oplus\Rr_M)$, then $g$ admits a
polyhomogeneous expansion $g \sim \sum_{j \geq 0} g_{m-j}$ with homogeneous components
corresponding to those of $q$, and $K=\Ff^{-1}[g]$ is Volterra (supported in $\{t \geq 0\}$)
and admits the associated asymptotic expansion as $t\downarrow 0^+$.
\label{Lem:Greiner}
\end{Lem}

\begin{proof}[Proof sketch]
For each fixed $(x,y)$ set $f(\xi,\tau) := q_x(y,\xi,\tau)$. Then $f$ is smooth on $(\A_x^{\ast}\times \Cc_-) \setminus \{0\}$, holomorphic in $\tau \in \Cc_-$, and parabolically homogeneous of degree $m$, i.e.\ $f(\lambda\xi, \lambda^2\tau) = \lambda^m f(\xi,\tau)$. Recall that $d = \mathrm{rk}\,\A_x$ and set $Q:=d+2$. Choose the minimal $j \in \Nn_0$ with $m+2j>-Q$.
Define the $j$-fold $\tau$-antiderivative
\[
f_j(\xi,\tau) := i^{j} \int_{\infty}^{\tau} \int_{\infty}^{\sigma_{j-1}} \cdots \int_{\infty}^{\sigma_1} f(\xi, \sigma_0)\, d\sigma_0 \cdots d\sigma_{j-1},
\]
where the integration runs along any piecewise $C^1$ path inside $\Cc_-$ and the constant of integration is fixed by requiring $f_j(\xi,\tau) \to 0$ as $|\tau| \to \infty$ within $\Cc_-$. Then $f_j$ is holomorphic in $\tau$, homogeneous of degree $m + 2j$, and $f_j \in L^1_{\mathrm{loc}}$ near $(\xi,\tau) = (0,0)$ because $m+2j > -Q$. Set
\[
g_x(y;\xi,\tau) := \partial_\tau^{j} f_j(\xi, \tau) \in S'(\A_x^\ast \times \Rr),
\]
which is a tempered distribution, homogeneous of degree $m$, and agrees with $f$ on $((\A_x^\ast\times\Cc_-) \setminus \{0\})$. By Paley-Wiener, $f_j$ being a boundary value from the open half-plane $\Cc_-$ implies $\F^{-1}_\tau[f_j](t)$ is supported in $\{t \geq 0\}$; applying $\partial_\tau^{j}$ corresponds to multiplication by $(it)^j$, hence preserves this support.
Equivalently, one may regularize by
\[
f_\varepsilon(\xi,\tau) := (1 + i \varepsilon \tau)^{-N} f(\xi,\tau), \ N \ \text{suff. large}, \ \varepsilon>0,
\]
so that $\supp \F^{-1}_\tau[f_\varepsilon] \subset \{t \geq 0\}$ for each $\varepsilon$, and $f_\varepsilon \to g_x(y;\cdot,\cdot)$ in $S'$ as $\varepsilon \downarrow 0$.
Let $K_x(y;\zeta,t): =\F^{-1}_{(\xi,\tau) \to (\zeta,t)}[g_x(y;\xi,\tau)]$. From the preceding paragraph, $K_x(y;\zeta,t) = 0$ for $t < 0$. This gives the claimed Volterra support after also taking the ordinary $\xi \leftrightarrow \zeta$ transform. Any two such extensions differ by a homogeneous distribution whose inverse $\tau$-Fourier transform is supported both in $\{t \geq 0\}$ and in $\{t \leq 0\}$, hence must vanish.
\end{proof}


\begin{Def}
A compactly supported anisotropic (possibly polyhomogeneous) operator $T \in \Psi_{ai, c}^{m}(\G \dtimes \Rr)$ is given by a family $T = (T_x)_{x \in M}$ such that 

\emph{i)} For all groupoid parametrizations $\Omega \cong U \times V$ with $U \subset M$ open and all $\kappa_1, \kappa_2 \in C_c^{\infty, 0}(\G \dtimes \Rr)$ with support in $\Omega$, the operator $\kappa_1 T \kappa_2$ is a compactly supported $C^{0}$-family of anisotropic (possibly polyhomogeneous) pseudodifferential operators of order $m$.

\emph{ii)} For each $\kappa_1, \kappa_2$ as above and with disjoint support, $\kappa_1 T \kappa_2 \in \Psi_{ai}^{-\infty}(\{x\} \times V)$. 
\label{Def:anisotropicOperator}
\end{Def}

Fix an algebroid connection $\nabla$ on $\A(\G)\oplus \Rr_M$ and denote the generalized exponential map by
\[
\Exp^\nabla \colon \A(\G)\oplus \Rr_M \longrightarrow \G\dtimes \Rr.
\]
Quantization is performed in exponential coordinates by the usual oscillatory integral with phase compatible with the parabolic dilations, yielding a map
\[
\Op_{ai} \colon S^{m}_{ai,ph}(\A^{\ast} \oplus \Rr_M)\ \longrightarrow\ \Psi^{m}_{ai,c}(\G \dtimes \Rr),
\]
well-defined modulo $\Psi^{-\infty}_{ai}$ and compatible with the $C^{\infty,0}$ structure.

Let $\mathcal{U} \subset \A\oplus \Rr_M$ be an open neighbourhood of the zero section
and $\mathcal{V} \subset \G \dtimes \Rr$ an open neighbourhood of the unit space such that
the generalized exponential map
\[
\Exp^{\nabla} \colon \mathcal{U} \iso \mathcal{V}
\]
is a diffeomorphism. Fix a groupoid parametrization centred at $x_0\in M$:
an open chart $\Omega \subset \G\dtimes \Rr$ with a diffeomorphism
\[
\psi \colon U\times V \iso \Omega \cap \mathcal{V},
\]
where $U \subset M$ and $V$ is an open neighbourhood of the unit in the longitudinal
coordinates of $\G_{x}$. Likewise choose an algebroid chart
\[
\theta \colon U \times W \iso \Omega \cap \mathcal{U},
\]
with $W$ an open neighbourhood of the zero section in $(\A\oplus \Rr_M)|_U$.
Define
\[
\alpha := \psi^{-1} \circ \Exp^{\nabla} \circ \theta \colon U \times W \longrightarrow U \times V.
\]
Then $\alpha(u,0) = (u,0)$, and for each $x = \varphi(u)$ the restriction
$\alpha_x \colon W_x\to V_x$ maps the fiber $(\A_x\oplus \Rr)\cap W$ diffeomorphically onto
$(\G_x\times \Rr) \cap V$ with $d\alpha_x|_{0} = \Id_{\A_x\oplus \Rr}$.
We have the commuting diagram
\[
\begin{tikzcd}[row sep=huge, column sep=huge, text height=2ex, text depth=0.5ex]
U\times W \arrow[d, "\theta"'] \arrow[r, "\alpha"] & U\times V \arrow[d, "\psi"] \\
\Omega\cap \mathcal U \arrow[r, "\Exp^\nabla"'] & \Omega\cap \mathcal V
\end{tikzcd}
\]

Write
\[
\Hom(\overline{r}^{\ast} E) := \overline{r}^{\ast} E\otimes (\overline{s}^{\ast}E)^{\ast} \longrightarrow \G \dtimes \Rr
\]
for the coefficient bundle of kernels. Shrinking $\Omega$ if necessary, we trivialize
$\Hom(\overline{r}^{\ast} E)$ over $\Omega \cap \mathcal{V}$ so that for each
$(\gamma,t) \in \Omega \cap \mathcal{V}$ there is a canonical identification
\[
\vartheta_{(\gamma,t)} \colon E_{\overline{r}(\gamma,t)} \otimes E_{\overline{s}(\gamma, t)}^{\ast}
\iso \End(E_x), \ x = \overline{s}(\gamma, t),
\]
e.g. by parallel transport along the base in $U$.

Using the charts $(\psi, \theta)$ and the trivialization $\vartheta$, we obtain a
canonical topological isomorphism 
\[
\Theta_{\psi, \theta} \colon 
\Psi_{ai}^{-\infty}\left(\overline{s}^{-1}(U) \cap \mathcal{V}, \overline{r}^{\ast} \Hom(E)\right)
\iso
\Psi_{ai}^{-\infty}\left(U \times W, \End(E_x)\right),
\]
defined by pulling back via $\psi^{-1}$, transferring $V$- to $W$-coordinates with $\alpha$,
and applying $\vartheta$ to identify coefficients with $\End(E_x)$, see also \cite{NWX}.

\begin{Prop}
Let $Q\in \Psi^{m}_{ai,c}(\G \dtimes \Rr, \overline{r}^{\ast} E)$ and let
$k_Q$ be its compactly supported Schwartz kernel (with values in
$\Hom(\overline{r}^{\ast} E)$). Then, shrinking $\Omega$ if necessary, there exists
a symbol
\[
q \in S^{m}_{ai}\left(U \times \Rr^{d+1}; \End(E_x)\right),
\ d := \mathrm{rk}\,\A_x,
\]
such that, in the local coordinates $(u,\zeta,t)\in U \times W$ (with $\zeta \in \A_x$),
\[
\Theta_{\psi,\theta}(k_Q)\ =\ k_q,
\]
where
\[
k_q(u;\zeta,t) = (2\pi)^{-(d+1)} \int_{\Rr^{d}} \int_{\Rr} e^{-i(\scal{\zeta}{\xi} + t \tau)} q(u,\xi,\tau)\, d\xi\, d\tau \in \End(E_x).
\]
Equivalently,
\[
k_{Q|_{\Omega \cap \mathcal{V}}} = \Theta_{\psi,\theta}^{-1}(k_q).
\]
The symbol $q$ is unique modulo $S^{-\infty}_{ai}$ and depends continuously on $Q$.
If $Q$ is polyhomogeneous, then $q$ is classical/polyhomogeneous in the anisotropic
sense.
\label{Prop:localizer}
\end{Prop}

Write
\[
\Sigma^{m,2}\left(\A^\ast \oplus \Rr_M; \End(E)\right) := S^{m}_{ai} / S^{m-1}_{ai}.
\]
An element $[q] \in \Sigma^{m,2}$ is represented by a fiberwise parabolically homogeneous, $\tau$-holomorphic function $q_x(y, \xi, \tau)$ of degree $m$, defined for $(\xi, \tau) \neq (0,0)$ and unique modulo $S^{m-1}_{ai}$. The multiplication on $\Sigma^{\bullet, 2}$ is pointwise composition in $\End(E)$: $[q_1] \cdot [q_2] := [q_1 q_2]$, well-defined since $S^{m-1}_{ai}$ is an ideal.
For $Q \in \widehat{\Psi}^{m}_{ai}(\G \dtimes \Rr; \overline r^\ast E)$, choose $q \in S^{m}_{ai,ph}$ with $Q = \Op_{ai}(q) \bmod \Psi^{-\infty}_{\G\dtimes\Rr, ai}$ and set
\[
\sigma_m(Q) := [q_m]\ \in \Sigma^{m,2},
\]
where $q_m$ denotes any leading homogeneous component of $q$; this is chart-independent and well-defined modulo $S^{m-1}_{ai}$. Moreover, $\sigma_{m+m'}(Q_1 Q_2)=\sigma_m(Q_1) \cdot \sigma_{m'}(Q_2)$.

\begin{Def}
A Volterra operator $Q \in \widehat{\Psi}_{ai}^m(\G \dtimes \Rr; \overline{r}^{\ast} E), \ m \in \Zz$ is a continuous operator $Q \colon C_c^{\infty}(\overline{r}^{\ast} E) \to C^{\infty}(\overline{r}^{\ast} E)$ such that
\begin{enumerate}
\item $Q$ has the Volterra property.
\item $Q = \Op_{ai}(q) + R$ for some symbol $q \in S_{ai}^m(\A^{\ast} \oplus \Rr_M, \overline{r}^{\ast} E)$ and some residual operator $R$ that is contained in the Vassout ideal with the Volterra property: $R \in \Psi_{\G \dtimes \Rr, ai}^{-\infty}(\G \dtimes \Rr)$.
\end{enumerate}

\label{Def:VolterraCalc}
\end{Def}

\begin{Rem}
Define the Payley-Wiener algebra as in \cite{LR} by $C_{PW}^{\infty}(\A^{\ast} \oplus \Rr_M) := \Ff C_c^{\infty}(\A \oplus \Rr_M)$. Given a cutoff $\chi \in C_c^{\infty}(\A \oplus \Rr_M)$ with $\chi \equiv 1$ near $(\zeta, t) = (0, 0)$, then $\widehat{\chi} := \Ff \chi \in C_{PW}^{\infty}$. We can rescale the kernel, preserving compact support $\chi_{\epsilon}(\zeta, t) := \epsilon^{-(d +2)} \chi(\epsilon^{-1} \zeta, \epsilon^{-2} t)$, $\chi_{\epsilon} \in C_{PW}^{\infty}$ and $q \mapsto \widehat{\chi}_{\epsilon} q$ furnishes a microlocal cutoff, leaving principal symbols unchanged. We set $C_{PW, +}^{\infty} := \Ff C_c^{\infty}(\A \oplus \Rr_M; t \geq 0)$ which is a Fr\'echet algebra with respect to pointwise composition, i.e. the Volterra property is preserved.
\label{Rem:PWalgebra}
\end{Rem}

\begin{Prop}
\emph{i)} The compactly supported calculus $\Psi_{ai,c}^{m}(\G \dtimes \Rr; \overline{r}^{\ast} E)$ and the extended calculus $\widehat{\Psi}_{ai}^m(\G \times \Rr; \overline{r}^{\ast} E)$ are closed with respect to composition that is given by groupoid convolution of reduced Schwartz kernels.

\emph{ii)} Let $P \in \widehat{\Psi}_{ai}^m(\G \dtimes \Rr; \overline{r}^{\ast} E)$ be such that its Volterra principal symbol is pointwise invertible (i.e. $\sigma_m(P)$ invertible in the symbol space $\Sigma^{m,2}$). Then there is $Q \in \widehat{\Psi}_{ai}^{-m}(\G \dtimes \Rr; \overline{r}^{\ast} E)$ such that $P Q = \Id - S_1, \ Q P = \Id - S_2$ with $S_1, S_2 \in \Psi_{\G \dtimes \Rr, ai}^{-\infty}(\overline{r}^{\ast} E)$. 
\label{Prop:VolterraCalc}
\end{Prop}

\begin{proof}
\emph{i)} Working in a chart provided by the localizer $\Theta_{\psi,\theta}$ by virtue of Prop. \ref{Prop:localizer}, let $Q_j=\Op_{ai}(q_j)$ with $q_j\in S^{m_j}_{ai,ph}(U \times \Rr^{d+1};\End(E_x))$ ($j=1,2$). Their reduced kernels are
\[
k_{q_j}(u; \zeta, t) = (2\pi)^{-(d+1)} \int e^{-i \scal{\zeta}{\xi} + t \tau)} q_j(u, \xi, \tau) \,d\xi\,d\tau,
\]
which vanish for $t<0$ by holomorphy in $\tau \in \Cc_-$. The composition kernel on $U$ is the
convolution $k_{q_1} \ast k_{q_2} = k_{q_1 \# q_2}$ is still supported in $\{t \geq 0\}$.
A standard computation shows that $q_1 \# q_2\in S^{m_1+m_2}_{ai,ph}$ with
\[
q_1 \# q_2\ \sim \sum_{\alpha} \frac{1}{\alpha!} \left(\partial_\xi^{\alpha} q_1 \right) \left(D_{\xi}^\alpha q_2\right),
\]
so $\Theta_{\psi,\theta}(Q_1 Q_2)  =\Op_{ai}(q_1\# q_2)$ modulo $\Psi^{-\infty}_{ai}$,
and the leading term gives $\sigma_{m_1+m_2}(Q_1 Q_2)=\sigma_{m_1}(Q_1) \cdot \sigma_{m_2}(Q_2)$.
\emph{ii)} Let $p := \sigma_m(P) \in \Sigma^{m,2}$ be pointwise invertible; write $p^{-1}$ for its
fibrewise inverse in $\Sigma^{-m,2}$. In a localizer chart, choose a representative
$p_m(u, \xi, \tau)$ of $p$ and define the first approximation $q_{-m}^{(0)}(u, \xi, \tau) := p_m(u,\xi, \tau)^{-1}\in S^{-m}_{ai}$. Define inductively $q_{-m-j}^{(0)} \in S^{-m-j}_{ai}$ so that, for each $N$,
\[
p\# \left(\sum_{j < N} q^{(0)}_{-m-j} \right) = 1 \ \bmod \ S^{-m-N}_{ai},
\]
with the anisotropic $\#$-product from \emph{(i)}.
This yields a formal inverse $q^{(0)} \sim \sum_{j \geq 0} q^{(0)}_{-m-j} \in S^{-m}_{ai,ph}$,
but the cutoffs used to globalize in $(\xi, \tau)$ need not preserve $\tau$-holomorphy,
so $q^{(0)}$ may fail to be Volterra. To make the kernel compactly supported and keep causality, pick $\chi \in C_c^\infty(\A \oplus \Rr_M)$ with $\chi \equiv 1$ near $(\zeta, t) = (0, 0)$ and $\supp \chi \subset \{t \geq 0\}$, and set $\widehat{\chi} := \F_{fib} \chi \in C_{PW,+}^\infty$. Replace $q$ by $\widehat{\chi} q$ and, for each homogeneous piece $q^{(0)}_{-m-j}$, apply the fibrewise Volterra extension given by Lem. \ref{Lem:Greiner} to get a unique homogeneous distribution $g_{-m-j}$ of degree $-m-j$ whose $(\xi, \tau)$-inverse Fourier transform is supported in $\{t \geq 0\}$, and which agrees with $q^{(0)}_{-m-j}$ off $(\xi, \tau) = (0,0)$. By a standard Borel summation as in \cite[p. 5-6]{PongeMika}, there exists
\[
q \in S^{-m}_{ai,ph} \ \text{such that} \
q \sim \sum_{j \geq 0} g_{-m-j}
\]
and $q$ is holomorphic in $\tau \in \Cc_-$; hence $\Op_{ai}(q)$ is Volterra.
 Define $Q:=\Op_{ai}(\widehat{\chi} q)$. By construction, $P Q = \Id - R_1$ and $Q P = \Id - R_2$ modulo $S^{-\infty}_{ai}$ in each chart.
\end{proof}

\section{Short time asymptotic expansion}


\subsection{Parabolic adiabatic deformation groupoid}

In order for us to specify the parabolic adiabatic deformation groupoid, we recall some of the necessary notions that enter into the filtered calculus as described in \cite{vEY}; see also \cite{DH, SLLS}. 
We fix the filtration $H^{\bullet}$ of maximal degree $N = 2$ on the Lie algebroid $\A \oplus \Rr_M \cong \A(\G \dtimes \Rr)$ of the groupoid $\G \dtimes \Rr$ that is given by $H^0 = \{0\}, \ H^1 = \A \oplus 0, \ H^2 = \A \oplus \Rr$. Fix the notation
\[
t^p H := H^p / H^{p-1}, \ \A(\G \dtimes \Rr)_H := \oplus_{i=0}^N H^i / H^{i-1} = \oplus_{i=0}^N t^i H. 
\]

Then $\A_{H, x} = \oplus_p t_x^p H$ has a canonical Lie algebra structure depending smoothly on the base point $x$. Let $\nabla$ denote a linear connection on $\A \oplus \Rr_M$ preserving the grading automorphism 
\[
\delta_{\lambda} \colon \A(\G \dtimes \Rr)_H \to \A(\G \dtimes \Rr)_H, \ \delta_{\lambda} = 
\oplus_{i=1}^{N} \lambda^{w_i}, \ w_1 = 1, \ w_2 = 2. 
\]
The $\delta_{\lambda,x}$ result from integration of the corresponding scaling automorphisms $\dot{\delta}_{\lambda, x}$ via the fiberwise exponential map 
\[
\exp^{\nabla} \colon \A(\G \dtimes \Rr)_H \to (\G \dtimes \Rr)_{H}; 
\]
the right hand side is a bundle of Lie groups and $\exp^{\nabla}$ implements a diffeomorphism of locally trivial bundles over $M$. Thereby, $\dot{\delta}_{\lambda, x} \in \Aut(t_x H)$ integrates to $\delta_{\lambda, x} \in \Aut((\G \dtimes \Rr)_{H, x})$. In particular 
\[
\delta_{\lambda} \colon (\G \dtimes \Rr) \to (\G \dtimes \Rr) 
\]
are bundle diffeomorphisms such that 
\[
\delta_{\lambda} \circ \exp = \exp \circ \dot{\delta}_{\lambda}. 
\]
Fix the splitting $\Psi \colon \A(\G \dtimes \Rr)_H \to \A \oplus \Rr_M$ corresponding to $\nabla$. This allows us to identify $\A(\G \dtimes \Rr)_H \cong \A \oplus \Rr_M$ via $-\Psi$ and with respect to the left-invariant vector fields on the bundles of Lie groups. Also, $(\G \dtimes \Rr)_{H, x} \cong \A_{H, x}$ via $\exp$ and with respect to the right-invariant vector fields on the Lie algebroid. We want to define the corresponding deformation Lie groupoid that is given as a set by the expression \eqref{DeformationGroupoid}. This groupoid is obtained by integration of the appropriate Lie algebroid of the deformation groupoid that is given by the set and the Lie algebra of vector fields:
\begin{align*}
& \U_{H^{\bullet}} := (\A \oplus \Rr_M) \times \Rr_{\hbar}^{\ast} \cup \A(\G \dtimes \Rr)_H \times \{0\}_{\hbar}, \\
& \Gamma(\U_{H^{\bullet}}) := \{X \in \Gamma((\A \oplus \Rr_M) \times \Rr_{\hbar}) : \partial_{\hbar}^i X \in \Gamma(H^i), \ i \geq 0\}. 
\end{align*}

We have the diagram
\[
\begin{tikzcd}[row sep=huge, column sep=huge, text height=2ex, text depth=0.5ex]
(\G \dtimes \Rr)_H \arrow{d} & \arrow[l, "\exp^{\nabla}"] \A(\G \dtimes \Rr)_H \arrow{d} \arrow[r, "-\Psi", "\simeq"'] & \A \oplus \Rr_M \arrow{d} \arrow{r} & M \times M \arrow[d, "\mathrm{pr}_1"] \\
M \ar[equal]{r} & M \ar[equal]{r} & M \ar[equal]{r} & M \arrow[bend right=60, swap]{u}{\Delta}
\end{tikzcd}
\]

Set $\nabla^{\Psi} := \Psi \circ \nabla \circ \Psi^{-1}$ for the induced connection on the filtered Lie algebroid $\A \oplus \Rr_M$ and 
\[
\Exp^{\nabla^{\Psi}} \colon \A \oplus \Rr_M \to \G \dtimes \Rr
\] 
the generalized exponential map. We have $\Gamma((\G \dtimes \Rr)^{pa}) \cong \Gamma(\A(\G \dtimes \Rr)_H \times \Rr_{\hbar})$, via $\widetilde{X} \mapsto ((x, \hbar) \mapsto \delta_{\hbar} \Psi^{-1} \widetilde{X}(x, \hbar))$ and the identification Lie algebra isomorphism 
\[
\Gamma(\A((\G \dtimes \Rr)^{pa})) = \{\widetilde{X} \in \Gamma(\A(\G \dtimes \Rr) \times \Rr_{\hbar}) : \partial_{\hbar|\hbar = 0}^k \widetilde{X} \in \Gamma(t^k H), \ k \in \Nn_0\}.
\]

The left hand side bracket is the bracked of $\A(\G \dtimes \Rr) \times \Rr_{\hbar}$ and the anchor is given by the anchor of $\A(\G \dtimes \Rr) \times \Rr_{\hbar}$. As in the definition of the standard (non-filtered) tangent groupoid, we indicate the smooth structure by charts of the form 
\[
(x, \xi, \tau, \hbar) \mapsto \begin{cases} \Exp_x^{\nabla^{\Psi}}(\Psi(\delta_{\hbar}(\xi, \tau), \hbar), \ \hbar \not= 0 \\
(x, \xi, \tau, 0), \ \hbar = 0 \end{cases}
\]

in a small tubular neighborhood of the zero section $O_M \subset \U \subset \A \oplus \Rr_M$. Denote by $(\G \dtimes \Rr)_H^{op}$ the groupoid with the opposite categorical structure. We have the evaluations
\[
\Gamma(\A((\G \dtimes \Rr)_H^{op})) \xleftarrow{\mathrm{ev}_0} \Gamma(\A((\G \dtimes \Rr)^{pa})) \xrightarrow{\mathrm{ev}_t} \A(\G \dtimes \Rr),
\]

and $\mathrm{ev}_0(\widetilde{X}) = \lim_{\hbar \to 0} \delta_{\hbar} \mathrm{ev}_{\hbar}(\widetilde{X})$. We have $\Gamma(\A(\G \dtimes \Rr)_H) = \Gamma(\A((\G \dtimes \Rr)_H^{op}))$ with respect to the left invariant vector fields of the Lie algebra on the right hand side, and the inclusions $(\G \dtimes \Rr)_{H, x} \hookrightarrow (\G \dtimes \Rr)_H$ are strict groupoid morphisms, $x \in M$. 

\subsection{Standard coordinates and lift}

We fix the notion of standard exponential coordinates on $(\G \dtimes \Rr)^{pa}$ that are obtained via a localizer associated to the generalized exponential map as described in the previous section. These are given by the $4$-tuple $(U, V, \varrho^{pa}, \U)$ where $U \times V \times \Rr_{\hbar} \supset \U \supset U \times \{0\} \times \{0\}_{\hbar}$ and $\U$ is invariant under the scaling action
\[
\alpha_{\lambda}(u, z, \hbar) = (u, \delta_{\lambda} z, \lambda^{-1} \hbar), \ (u, z, \hbar) \in \U, \ \lambda > 0. 
\]

The map $\varrho$ is the specialization of the Lie algebroid anchor of $\A \oplus \Rr_M$ with respect to the algebroid parametrization; see also \cite{McDonald}. 


Let $\{\mu_x\}_{x \in M}$ be a Haar system for $\G$, and endow $\G \dtimes \Rr$ with the product Haar system $\{\mu_x \otimes dt\}_{x \in M}$. On $(\G \dtimes \Rr)^{pa}$ we use the following specialization:
\begin{itemize}
\item For $\hbar \neq 0$, in the chart $U \times V$ with coordinates $(x, z)$, we pull back the Haar density by
\begin{align*}
& \Exp_x^{\nabla^{\Psi}} \circ \delta_{\hbar} \colon V \to (\G \dtimes \Rr)_{x}, \\
& \delta_{\hbar}(z) := (\hbar \zeta, \hbar^2 t), \ \text{if} \ z = (\zeta, t),
\end{align*}

and write
\begin{align*}
& d\mu_{x, \hbar}(z) = J_x(z, \hbar)\,dz, \
J_x(z, \hbar) := \left|\det D \left(\Exp_x^{\nabla{\Psi}} \circ \delta_{\hbar}\right) \right|.
\end{align*}

\item For $\hbar = 0$, on the osculating group $(\G \dtimes \Rr)_{H, x} \cong \A_x \oplus \Rr$ we take the left Haar measure given by the product of a fixed smooth density on $\A_x$ and Lebegue measure $dt$, so in the $(\zeta, t)$-coordinates
\[
d\mu_{x, 0}(\zeta, t) = d\zeta \, dt.
\]
\end{itemize}

These choices are compatible with parabolic scaling:
\[
(\delta_{\lambda})^{\ast}(d \mu_{x, \hbar}) = \lambda^{d+2} \,d\mu_{x, \lambda \hbar}, \ \lambda > 0
\]

and glue to a smooth Haar system $\{\mu_{x, \hbar}\}_{(x, \hbar) \in M \times \Rr_{\hbar}}$ on $(\G \dtimes \Rr)^{pa}$. 

Let $q \in S_{ai, ph}^m(\A^{\ast} \oplus \Rr_M; \End(E))$ be a Volterra symbol for $T$ in the local chart. Define the \emph{$\hbar$-rescaled symbol}
\begin{align}
q_{\hbar}(u, \xi, \tau) := q\left(u, \hbar \xi, \hbar^2 \tau\right), \ \hbar \geq 0, \label{paSymbol}
\end{align}

where for $\hbar = 0$ the right-hand side is understood as the (homogeneous) principal limit on $(\xi, \tau) \not= (0, 0)$ (i.e. the Volterra principal symbol in degree $m$ on the osculating fibers). Then the quantization on $(\G \dtimes \Rr)^{pa}$ is given, in standard coordinates, by the fiberwise oscillatory integral
\begin{align}
k_q^{pa}(u, \zeta, t, \hbar) &:= (2 \pi)^{-(d+1)} \int_{\Rr^d} \int_{\Rr} e^{-i \scal{\zeta}{\xi} + i t \tau} q_{\hbar}(u, \xi, \tau) \,d\xi\,d\tau \label{paQuant}
\end{align}

which defines a family of reduced kernels supported in $\{t \geq 0\}$. Equivalently, on the kernel side this is just the parabolic rescaling
\begin{align}
& k_q^{pa}(u, \zeta, t, \hbar) = \hbar^{-d-2} k_q(u, \zeta \hbar, t \hbar^2), \ \hbar > 0 \notag \\
& k_q^{pa}(u, \zeta, t, 0) = k_{q_m}(u, \zeta, t), \label{paRescale}
\end{align}

where $k_q$ is the reduced kernel of $T = \Op_{ai}(q)$ and $q_m$ is the leading homogeneous component of $q$. 

We continue to work in the standard coordinates, i.e. with respect to a groupoid parametrization of $(\G \dtimes \Rr)^{pa}$ that is refined via the generalized exponential map $\Exp^{\nabla^{\Psi}}$ with respect to a splitting $\Psi$. 
In these standard coordinates, we study compactly supported continuous families $f \in C(U \times \Rr_{\hbar}, C_c^{\infty}(V))$. The goal is to extend a given Volterra operator $T$ to a $1$-parameter family on the parabolic adiabatic deformation grouopoid $(\G \dtimes \Rr)^{pa}$. The family $\{T_{\hbar}\}_{\hbar > 0}$ acts, for $s$-fibered (longitudinally smooth, transversely continuous) distributional kernels $k \in \mathscr{E}_{s}^{', 0}$ via
\begin{align}
(T_{\hbar} g)(u) &= \int_V k_q^{pa}(u, z, \hbar) g(\exp(\varrho^{pa} \delta_{\hbar} z) u) \,dz \notag \\
&= \hbar^{-d-2} \int_V k_q^{pa}(u, \delta_{\hbar}^{-1} z, \hbar) g(\exp(\varrho^{pa} z) u)\,dz, \label{paVolterra}
\end{align}

and for $\hbar = 0$ it is the model operator on the osculating bundle
\[
(T_0 g)(u) = \int_{\A_x \times \Rr} k_{q}^{pa}(u, \zeta, t, 0) g\left(\exp(\varrho^{pa}(\zeta, t))u \right)\, d\zeta\,dt.
\]
Recall that by an application of the scaling action (cf. \cite{DS}) where we write $k = k^{pa}$
\[
((\alpha_{\lambda})_{\ast} k, \varphi)(u, \hbar) = \lambda^{d+2}(k, \varphi \circ \alpha_{\lambda})(u, \lambda \hbar), \ (u, \hbar) \in r(\U),
\]

where the $s$-fiberwise action of $k$ is given by
\[
(k, g)(u, \hbar) = \int_V k(u, z, \hbar) g(u, z, \hbar) \,dz, \ g \in C^{\infty, 0}(\U). 
\]

The extended Volterra operator family kernel $k$ then fulfills the homogeneity property
\begin{align}
& \forall_{\lambda > 0} \ \lambda^{-m-d-2} (\alpha_{\lambda})_{\ast} k - k \in C^{\infty, 0}(\U). \label{Homogeneity}
\end{align}

We make use of these expressions in standard coordinates to study the asymptotic expansion of the heat kernel on a Lie groupoid. 

\begin{Prop}
Let $A \in \Diff^m(\G; r^{\ast} E)$ be a $\G$-differential operator acting on sections of a vector bundle $\pi \colon E \to M$ on the $d$-dimensional compact manifold (possibly with corners) $M$ of order $m > 0$ which is self-adjoint and elliptic, acting as an unbounded operator on the $C_r^{\ast}(\G)$-module $\E := C_r^{\ast}(\G, r^{\ast} E)$ and non-negative with respect to the canonical $L^2 = H^0$ inner product. Then the semigroup heat kernel $k_t$ of $\exp(-tA)$ has a short time asymptotic expansion when restricted to the diagonal $\Delta_M$ in $\G$ of the form
\[
k_{t|\Delta} \sim \sum_{j=0}^{\infty} t^{\frac{j-d}{m}} q_j, \ \text{as} \ t \to 0^{+}. 
\]

Here the coefficients are sections $q_j\in \Gamma\left(\End(E)\otimes |\!\wedge|\A^\ast\right)$ that depend continuously on $x \in M$. 
\label{Prop:AsymptExp}
\end{Prop}

\begin{proof}
Denote by $C_+(\Rr, H^0)$ the set of continuous functions $\psi \colon \Rr \to H^0(r^{\ast} E)$ such that there is a $t_0 > 0$ with $\psi(t) = 0$ for $t \leq t_0$, cf. \cite{Greiner}. Define $Q \colon C_+(\Rr, H^0) \to C_+(\Rr, H^0)$ via 
\[
(Q \psi)(s) := \int_{-\infty}^0 e^{-(s+t) A} \psi(t)\,dt, 
\]

where we utilize the earlier defined heat semigroup. By the strong continuity of the semigroup, Prop. \ref{Prop:HeatSemigroup}, we have $Q \psi \in C_+(\Rr, H^0)$. The $H^0$-contractivity of the semigroup furnishes
\[
\|(Q \psi)(s)\| \leq |s - t_0| \sup_{[t_0, s]} \|\psi(t)\|, \ \forall \ \psi \in C_+ 
\]

such that $\supp \psi \subset [t_0, \infty)$. We have
\[
C_+(\Rr, H^0) = \varprojlim_{t_0} C_{[t_0, \infty)}(\Rr, H^0)
\]

with the projective limit topology of uniform convergence on compact subsets over $t_0 \in \Rr$. We have the continuous inclusions $C_c^{-\infty, 0} \hookrightarrow C_+ \hookrightarrow C_c^{\infty, 0}$ and hence $Q$ furnishes a $C^{\infty, 0}$ Schwartz kernel $k_Q$ that is contined in $\Gamma^{0, \infty}(r^{\ast} E \otimes s^{\ast} E^{\ast})$. Let the heat operator corresponding to $A + \partial_t$ be denoted by $\widetilde{Q}$, i.e. the natural realization of the fundamental solution, with fixed initial condition, inside the extended Volterra calculus $\widehat{\Psi}_{ai, ph}^{-m}$. Denote by $\widetilde{Q}_{\hbar}$ its lift to a $1$-parameter family on the parabolic adiabatic groupoid. Denote by $\mathcal{P}^s$ the set of $C^{\infty}$ families of polynomial volume densities of homogeneous degrees on the fibers of $(\G \dtimes \Rr)_{H, x} \cong \A(\G \dtimes \Rr)_{H, x}$ defined by 
\[
\mathcal{P}^s((\G \dtimes \Rr)_H) := \{k \in \Psi_{ai}^{-\infty}((\G \dtimes \Rr)_H; \widetilde{r}^{\ast} E)|(\delta_{\lambda})_{\ast} k = \lambda^s k, \ \forall \ \lambda > 0\}. 
\]

Express the lifted family $\widetilde{Q}_{\hbar}$ in fixed standard coordinates $\U$. Then the task is to study the asymptotic behavior of the localizer $\Theta^{\ast}(k_{\widetilde{Q}_{\hbar}|\U})$ at $\hbar = 0$ as $t \to 0^{+}$. 
%
%
We have
\[
\Theta^{\ast}(k_{\widetilde{Q}_{\hbar}|\U}) - \sum_{j=0}^{J_N} q_j \in C^N
\]

and by \cite[Lem. 2]{DH}
\begin{align}
& (\delta_{\lambda}^{(\G \dtimes \Rr)_H})_{\ast} q_j = \lambda^{-m-d-2-j} q_j - \lambda^{-m-d-2-j} \log |\lambda| p_j, \label{Homogeneity}
\end{align}

where $p_j$ is $C^{\infty}$ and strictly homogeneous, i.e. 
\[
p_j \in \mathcal{P}^{-m-j-d}((\G \dtimes \Rr)_H; \End(E))
\]

such that
\[
(\delta_{\lambda}^{(\G \dtimes \Rr)_H)})_{\ast} p_j = \lambda^{-m-d-j} p_j, \ \lambda \not= 0.
\]

Here the representatives $q_j$ satisfying \eqref{Homogeneity} are unique mod $\mathcal{P}^{-m-d-j}((\G \dtimes \Rr)_H; \End(E))$. We have
\[
(\Id + R_1) Q \psi = \widetilde{Q} \psi = Q (\Id + R_2) \psi
\]

where the $R_j$'s are contained in the Vassout ideal $\Psi_{\G \dtimes \Rr, ai}^{-\infty}$. By virtue of this identity we obtain
\[
\Theta^{\ast}(k_{Q_{\hbar}|\U}) \sim \sum_{j=0}^{\infty} q_j.
\]

By virtue of Taylor's theorem, by adding terms in $\mathcal{P}^{-m-d-j}$ to $q_j$, we have that for every $N$ there exists $j_N$ such that
\[
\left(\Theta^{\ast}(k_{Q_{\hbar = 0|\U}}) - \sum_{j=0}^{j_N} q_j\right)(\widetilde{\zeta}) = O(|\widetilde{\zeta}|^N)
\]

as $\widetilde{\zeta} \to \widetilde{o}_x \in (\G \dtimes \Rr)_{H, x}$ uniformly for $x$ in compact subsets of $M$. 
Take the fiberwise Fourier transform in $(\zeta, t)\mapsto (\xi, \tau)$:
\[
\widehat{q}_j(\xi, \tau):=\F_{(\zeta, t)\to (\xi, \tau)}[q_j],
\widehat{p}_j:=\F[p_j].
\]
Because $\supp q_j \subset \{t \geq 0\}$, the Paley-Wiener theorem implies that for each
$\xi \in \Rr^d$, the map $\tau \mapsto \widehat{q}_j(\xi,\tau)$ is the boundary value (from
$\Im\tau < 0$) of a holomorphic function with at most polynomial growth. Since $p_j$ is a
polynomial density in $(\zeta, t)$, $\widehat{p}_j$ is a finite linear combination of
derivatives of the Dirac mass supported at $(\xi, \tau) = (0, 0)$.

Fourier transform intertwines dilation pushforwards by
\[
\widehat{(\delta_\lambda)_{\ast} u}(\xi, \tau) = \lambda^{-(d+2)} \widehat{u}(\lambda^{-1} \xi,\lambda^{-m} \tau),
\]
hence \eqref{Homogeneity} becomes
\begin{equation}\label{FT-homog}
\widehat{q}_j(\lambda^{-1} \xi, \lambda^{-m} \tau) = \lambda^{-(m+j)} \widehat{q}_j(\xi, \tau)
- \lambda^{-(m+j)} \log(\lambda) \widehat{p}_j(\xi, \tau),\ \lambda>0,
\end{equation}
as an identity of tempered distributions on $\Rr_{\xi}^d \times \Rr_{\tau}$.
Since $\widehat{p}_j$ is supported at $(\xi, \tau) = (0, 0)$, \eqref{FT-homog} reduces on
$(\Rr^d \times \Rr) \setminus \{(0,0)\}$ to
\[
\widehat{q}_j(\lambda^{-1}\xi, \lambda^{-m}\tau) = \lambda^{-(m+j)} \widehat{q}_j(\xi, \tau),\ \lambda > 0,
\]
so $\widehat{q}_j$ is strictly homogeneous of degree $m+j$ there.
Fix $\phi \in S(\Rr_{\xi}^{d})$ and test \eqref{FT-homog} against $\phi(\xi) \eta(\tau)$
with $\eta \in S(\Rr_\tau)$ arbitrary. Define the $\tau$-distributions
\[
H(\eta) :=\scal{\widehat{q}_j(\xi,\cdot)}{\phi(\xi)\otimes \eta}, \
P(\eta) := \scal{\widehat{p}_j(\xi,\cdot)}{\phi(\xi)\otimes \eta}.
\]
By Paley-Wiener, $H$ is the boundary value on $\Rr$ of a holomorphic function on the lower
half-plane $\Cc_-$ with polynomial growth. In contrast, $P$ is a finite linear combination of
derivatives of $\delta$ at $\tau = 0$ since $\widehat{p}_j$ is supported at $\xi=0$ and
$\tau=0$, hence $\supp P\subset\{0\}$.
Now \eqref{FT-homog} yields, for all $\lambda>0$,
\[
H(\eta_{\lambda}) = \lambda^{-(m+j)} H(\eta) - \lambda^{-(m+j)} \log(\lambda) P(\eta), \
\eta_{\lambda}(\tau) := \int_{\Rr^d} \phi(\xi) \eta(\lambda^{-m}\tau)\,d\xi,
\]
and in particular we can fix $\phi$ with $\int \phi = 1$ and absorb it in the notation so that
$\eta_{\lambda}(\tau) = \eta(\lambda^{-m}\tau)$. For any $\tau_0$ with $\Im \tau_0 < 0$ choose $\eta$ supported in a small neighbourhood of $\tau_0$. Then $P(\eta) = 0$ since $\supp P\subset\{0\}$, and we obtain the identity
\[
H(\eta_{\lambda}) = \lambda^{-(m+j)} H(\eta), \ \Im \tau_0 < 0.
\]
By uniqueness of analytic continuation, this identity holds as an equality of boundary values for all
real $\tau$ as well, and hence the term proportional to $\log \lambda$ must vanish:
$P(\eta)=0$ for all $\eta \in S(\Rr)$. Therefore $P=0$ as a distribution, hence $\widehat{p}_j = 0$ and thus $p_j=0$.
Expressing everything in the localizer with respect to standard coordinates, we find:
\[
k_Q(x, s-t) = \sum_{j=0}^{j_N} \Theta(\check{q}_{j,x}(o_x, t-s)) + O(|t|^{\frac{N}{r}})
\]

as $|t| \to 0$ uniformly in $x$ and, making use of the homogeneous dimension $m + d + 2$ of $M \times \Rr$
\[
\Theta_x(\check{q}_{j,x}(o_x, t)) = t^{\frac{j-d-2}{m}} \Theta_x(\check{q}_x(o_x, 1))
\]

for each $x, t > 0$ where we made use of the left-trivialization of $(\G \dtimes \Rr)_{H, x}$ to identify the two sides. Setting $\check{q}_j(x) \,dt := \Theta_x(\check{q}_{j,x}(o_x, 1))$ furnishes the result.
\end{proof}

\end{document}